OpenStudy (anonymous):

Which statement is true? In any right triangle, the sine of one acute angle is equal to the cosine of the other acute angle. In any right triangle, the sine of one acute angle is equal to the sine of its complementary angle. In any right triangle, the cosine of one acute angle is equal to the cosine of its complementary angle. In any right triangle, the sum of the sine of one acute angle and the cosine of the other acute angle is 1.

OpenStudy (anonymous):

D i think

OpenStudy (anonymous):

yea thanks man

OpenStudy (wolf1728):

No I don't think D sine 45 = .707 cosine 45 = .707 Sum does NOT = 1

OpenStudy (wolf1728):

sine 10 = .17365 cosine 80 = .17365 sum NOT = 1

OpenStudy (wolf1728):

Seems like it might be A Not sure

OpenStudy (wolf1728):

Anyone else care to give an answer?

OpenStudy (anonymous):

@wolf1728 is correct, it's A. Those are cofunction identities. \[\cos(90^o-\theta)=\sin \theta\] \[\sin(90^o-\theta)=\cos \theta\]